# AtomSurf: Surface Representation for Learning on Protein Structures

### v3

@@ Abstract
We demonstrate that the resulting architecture achieves state-of-the-art results on all tasks in the Atom3D benchmark, while adhering to the strict benchmark protocol, as well as more broadly on binding site identification and binding pocket classification. Furthermore, we use coarsened surfaces and optimize our approach for efficiency, making our tool competitive in training and inference time with existing techniques. Our code and data can be found online: https://github.com/Vincentx15/atomsurf
@@ 1 Introduction
Structural bioinformatics data is becoming available at an unprecedented pace. Advances in cryogenic Electron Microscopy (cryo-EM) in particular, have led to the production of evermore experimentally derived structures, as well as larger systems and better resolutions Fontana et al. (2022). The development of AlphaFold Jumper et al. (2021) along with many subsequent works have made protein structures abundantly available, with over a million high-quality predictions in the Protein Data Bank (PDB) Berman (2000) and over 600 million in the ESM Metagenomic Atlas (ESMatlas) Lin et al. (2022). There is thus a growing demand for machine learning techniques which can leverage this structural data to help advance the fields of structural bioinformatics and drug design.
Protein structures are complex objects characterized both by atomic coordinates as well as intricate bio-chemical interactions between them that depend on their geometry. To be used in a learning pipeline, an initial modeling step transforming protein structures into a well-defined mathematical object is necessary. Different mathematical representations encode different structural and biological priors. For instance, the point cloud representation disregards the connectivity induced by chemical interactions, but allows for the most generic geometric description of the data. Protein surfaces choose to trade fine-grained information of the interior of a protein for an accurate depiction of its outer surface. This representation is thought to be of particular interest to study active interaction sites that mostly depend on properties of the surface because of the screening effect. However, even for interactions dominated by surface terms, the knowledge of the interior can encode the stability of the surface. These different representations are illustrated in Figure 1.
Figure 1: Diverse mathematical objects used to represent a protein structure, sequences, molecular surfaces (blue), atom-level and residue-level point clouds (red) and graphs (green). Effective machine learning for protein structures hinges on selecting the appropriate mathematical representation along by a compatible machine-learning technique.
@@
Beyond using a single representation for proteins, the simultaneous representation of a protein as several mathematical objects holds promise. Indeed, different representations encode different biological priors of the data and present different computational advantages. A well-studied combination is the use of sequence information along with a graph representation of the structure. For instance, in Hermosilla et al. (2020); Fan et al. (2022) the authors enrich the graph with additional edge types that encode the sequence. Another way to include sequence information in a graph, is to use sequence embeddings, especially ones derived from protein language models and hence benefiting from the large amounts of available sequence data (Wu et al., 2023; Zhang et al., 2023).
@@ 2 Motivation and Contribution
Despite this recent progress in comparing different representations for learning on protein data, relatively less focus has been given to surface-based representations, even though they have shown promising results in several applications (Gainza et al., 2020; Sverrisson et al., 2021; Wang et al., 2023). Approaches based on the surface representation have typically followed the initial MaSIF paper validation (Gainza et al., 2020), and hence have never been directly compared to other representations in the context of a single well-established benchmark. At the same time, powerful surface-based encoders have recently been proposed in the geometry processing/computer graphics literature, such as DiffusionNet (Sharp et al., 2022) significantly outperforming, in terms of both robustness and accuracy, the early Geodesic-CNN based techniques Masci et al. (2015); Monti et al. (2017) which formed the basis of (Gainza et al., 2020). Unfortunately, it is not currently known how the best currently available surface-based encoders compare to other learning-based paradigms in the protein analysis tasks (e.g., on the Atom3d benchmark).
@@
  • State-of-the-art results in a wide variety of challenging scenarios.
  • Enhanced computational throughput by using residue graphs and coarsened meshes.
The rest of the paper is organized as follows: in Section 3.1, we present the surface representation we use and the specialized networks used to process it. Section 3.2 highlights a challenge for surface networks learning on various scales and presents solutions to mitigate these issues. In Section 3.3, we propose to synergistically integrate graph and surface information within a unified architecture, harnessing the power of both representations. We provide the details regarding the chosen architecture in Section 3.4 and analyze its computational aspects in Section 3.5.
@@ 3.1 Surface Representation Learning
Our first objective is to study the utility of the best existing surface encoders for learning on protein data. To generate the surface representation $\mathcal{S}_{\mathbf{P}}$ of the protein $\mathbf{P}$, our initial step involves computing the protein surface using MSMS (Sanner et al., 1996). The resulting meshes are usually quite large, but in the rare cases of small proteins, we incrementally increase sampling density until we achieve a minimum number of 256 vertices. Then, we employ quadratic decimation (Garland & Heckbert, 1997), to embed our surfaces into coarser and more compact meshes, again ensuring a minimum vertex count. Coarsening the mesh has an impact on computational efficiency and on performance, assessed in Sections 3.5 and 4. Finally, we ensure mesh quality by removing non-manifold edges, duplicated or degenerate vertices and faces, as well as small disconnected components. We compute several surface features: the Gaussian and mean curvature, as well as the shape index at each vertex. We also include the normal as well as the heat kernel signature (Sun et al., 2009) at each point.
At the basis for our surface-based learning method, we employ the DiffusionNet approach (Sharp et al., 2022). This method has been proven to be highly robust and effective across a diverse set shape analysis tasks (e.g., Attaiki et al. (2021); Cao & Bernard (2022); Sun et al. (2023); Li et al. (2022) among others). In particular, DiffusionNet avoids using local patch parametrizations, which can lead to instabilities across different mesh structures and enables long range and multi-scale information propagation by using learned diffusion for information sharing. The mathematical foundation of DiffusionNet is the heat equation, which simulates the diffusion of heat, or, equivalently, the behavior of Brownian motion on a surface over time. For a surface $S$, let $\mathbf{f}_{t}:S\to\mathbb{R}$ be the function that defines the heat distribution over $S$ at time $t$ and $\Delta_{S}$ the Laplace-Beltrami operator (Meyer et al., 2003; Vallet & Levy, 2008) of the surface. The heat equation is the linear differential equation:
$\frac{\partial\mathbf{f}}{\partial t}$ $=\Delta_{S}\mathbf{f}.$ (1)
Its solution is obtained by diagonalizing the Laplace-Beltrami. In practice, we store the $k=128$ smallest eigenvalues of $\Delta_{S}$ in a diagonal matrix $\Lambda\in\mathbb{R}^{k\times k}$, with the corresponding eigenvectors in $\Phi\in\mathbb{R}^{n\times k}$, and vertex area weights $M\in\mathbb{R}^{n\times n}$. The spectrally truncated solution to the heat equation is given as $\mathbf{f}_{t}=\Phi e^{-\Lambda t}(\Phi^{T}M)\mathbf{f}_{0}$.
In DiffusionNet, this equation is used to perform information propagation on a surface using a feature map as $\mathbf{f}_{0}$ and learning the diffusion time in a task-specific manner. This mechanism relies on dense linear algebra operations, offering straightforward differentiation with respect to both $\mathbf{f}$ and $t$. The diffusion layers are combined with features based on $\mathbf{f}_{t}$, their spatial gradients and standard, point-wise MLPs. This leads to an architecture, which can capture multi-scale geometric details of a surface in a task-specific manner. To establish a benchmark method for surface analysis on Atom3d, we utilize three DiffusionNet blocks, applying them to the protein surfaces $\mathcal{S}_{\mathbf{P}}$ under consideration.
@@ 3.2 Adapting to Protein Surfaces of Diverse Scales
To address this issue, we enhance the original DiffusionNet framework in two ways. First, we enable support for batch (the original model was limited to batch sizes of one) and incorporate a Batch Normalization layer Ioffe & Szegedy (2015) after each diffusion layer to stabilize learning. Second, we facilitate the optimization process by incorporating biological priors relevant to spatial scales. Consequently, we determined that diffusion times around 10 resulted in receptive fields around 10 Å (see Supplementary Figure 6), which aligns with the spatial scale of binding sites. Inspired by the inherent multi-scale nature of protein structures, we opted to draw samples from a normal distribution $t\sim\mathcal{N}(10,5)$, characterized by a relatively high variance. The absolute values of these samples were then utilized as the initial values for our diffusion timescales. Both the large scale and the large variance are retained during training, as illustrated in Figure 7), enabling efficient multi-scale and long-distance message passing. These enhancements to DiffusionNet implementation mitigate the instabilities in the training process (as seen in Figure 13) and are available in the provided code repository and as a pip package.
@@ 3.3 Hybrid Representation Learning
As mentioned earlier, beyond assessing the efficacy of surface-based learning compared to other representations for protein learning, we explore the benefits of integrating different representations in a unified framework, harnessing their distinct strengths. Intuitively, surface representations can capture the intricate geometric details critical for tasks involving protein interactions, while graph representations detail the specific atomic interactions within a protein’s interior that indirectly influence its surface dynamics and interaction capabilities. Furthermore, these representations facilitate complementary approaches to learning: local message passing through graphs and global information dissemination for surfaces via learned diffusion. Inspired by these considerations, we propose a method that enables feature sharing between graph and surface-based representations. As emphasized below, unlike previous related approaches Somnath et al. (2021); Pegoraro et al. (2024), we enable communication between the two representations across all learned layers of the network.
As a basis for our hybrid representation, in addition to the surface $\mathcal{S}_{\mathbf{P}}$, we construct a graph representation $\mathcal{G}_{\mathbf{P}}=(\mathcal{V}_{g},\mathcal{E}_{g})$. To ensure consistency and fairness in comparison within the Atom3d benchmark, we follow their conventions: each node in the graph corresponds to an atom within the protein and edges $\mathcal{E}_{g}$ are defined between pairs of atoms that are within a 4.5 angstrom radius cutoff. The only initial node features are one-hot encoding of atom types.
Outside of the benchmark setting, we add edge features to our bipartite graph. Given an edge $e_{g,s}$ tying a graph node to a surface vertex with normal $\vec{n}_{s}$, we use its direction $\vec{u}_{g,s}=\frac{x_{s}-x_{g}}{||x_{s}-x_{g}||}$ as a vector edge feature. The associated distance and angle, $\langle\vec{n}_{s},\vec{u}_{g,s}\rangle$ are encoded using sixteen radial basis Gaussian functions, and used as scalar edge features in the bipartite graph. We follow a similar protocol for edges going from a vertex to a graph node.
@@
We now define block operations to encode a protein using $\mathcal{S}_{\mathbf{P}},\mathcal{G}_{\mathbf{P}}$ and $G$. Denote encoders on surfaces and graphs as $s_{\theta}$ and $g_{\theta}$, respectively, and the set of input features as $\mathcal{X}=\{x_{n},n\in\mathcal{V}\}$. The corresponding encoded features are =$\mathcal{H}=\{h_{n},n\in\mathcal{V}\}$ with $h_{n}=s_{\theta}(x_{n})$ for nodes $n\in\mathcal{V}s$ and $h_{n}=g_{\theta}(x_{n})$ for nodes $n\in\mathcal{V}_{g}$. Our general methodology incorporates message-passing neural networks, denoted $\text{MP}_{\theta}$, over the bipartite graph $G$, such that at a layer $l$, we get $\mathcal{X}^{l+1}=\text{MP}^{l}_{\theta}(\mathcal{H}^{l})$. By employing distinct sets, $\theta_{sg}$ and $\theta_{gs}$, the architecture handles messages traversing from the surface to the graph and vice versa. Those block operations can be stacked as shown in Figure 2. We emphasize that our feature sharing occurs on the local (node) level and is enabled by the proximity relations in 3D space. Moreover, our hybrid approach trains both representations jointly, while enabling information sharing across all network layers, which is crucial to its success.
@@ 3.4 Proposed architectures
Different bipartite message passing networks and organization are possible, encompassing as special cases several existing approaches. We provide the implementation details in Appendix C. We evaluate these configurations and conduct several ablation studies on our final model, in Section 4.6.
@@ 3.5 Computational enhancements
Surface-methods are traditionally thought to be compute-expensive methods, motivating approaches to side-step their intrinsic complexity (Sverrisson et al., 2021). Through a complexity analysis presented in Section D.1, we found the number of vertices to be critical in DiffusionNet runtime, which we addressed by coarsening our meshes. In this coarse surface regimen, the graph encoding with ProNet is the computational bottleneck, which is smaller than when used in isolation. Hence, we claim that our method is faster than this efficient graph encoder in the coarse surface regimen.
From the storage and memory perspectives however, the memory footprint of the surface-related operations dominate the ones originating from the graphs, especially because of stochasticity. We did not find it to be limiting in terms of I/O, but rather in terms of batch size. We implemented a dynamic batching procedure, alleviating our memory issues. Further details are provided in D.2. Finding ways to reduce the memory footprint of surface networks remains an important direction.
@@ 4.1 Performance of Surface Representation
We start by validating our surface encoder on the RNA segmentation benchmark (Poulenard et al., 2019) for surface methods. The task is to segment 5s ribosomal RNA molecules into functional components. We assess the impact of the proposed enhancements to DiffusionNet by showing the learning curves of the enhanced models on the RNA segmentation task (see Figure 13). Moreover, we compare their performance to other recent surface encoders, DGCNN (Wang et al., 2019) and DeltaConv (Wiersma et al., 2022) and report results in Figure 13. Our experiments show that our adjustments significantly improve the DiffusionNet stability issues, and allow it to converge to a model with a test accuracy of 84.1% instead of of 80.9%. This accuracy is significantly higher than DGCNN (74.7%) and DeltaConv (78%).
Then, to compare the performance of our surface encoder to the use of other representations, we turn to the Atom3d benchmark, focusing on its three tasks exclusive to proteins. The Protein Interaction Prediction (PIP) task aims to predict which part of a protein interacts with another, and holds about 120k examples. The Mutation Stability Prediction (MSP) task is to determine if a mutation enhances the stability of protein-protein interaction, with about 5k examples. Finally, the Protein Structure Ranking (PSR) task aims to assign a quality score to predicted protein structures with around 10k proteins. A more detailed description of our datasets is available in Appendix E
@@
Auroc Auroc Local R Global R

### v4

@@ Abstract
We demonstrate that the resulting architecture achieves state-of-the-art results on all tasks in the Atom3D benchmark, while adhering to the strict benchmark protocol, as well as more broadly on binding site identification and binding pocket classification. Furthermore, we use coarsened surfaces and optimize our approach for efficiency, making our tool competitive in training and inference time with existing techniques. Code can be found online: github.com/Vincentx15/atomsurf
@@ 1 Introduction
Structural bioinformatics data is becoming available at an unprecedented pace. Advances in cryogenic Electron Microscopy (cryo-EM) in particular, have led to the production of evermore experimentally derived structures, as well as larger systems and better resolutions (Fontana et al., 2022). The development of AlphaFold (Jumper et al., 2021) along with many subsequent works have made protein structures abundantly available, with over a million high-quality predictions in the Protein Data Bank (PDB) (Berman, 2000) and over 600 million in the ESM Metagenomic Atlas (ESMatlas) (Lin et al., 2022). There is thus a growing demand for machine learning techniques which can leverage this structural data to help advance the fields of structural bioinformatics and drug design.
Protein structures are complex objects characterized both by atomic coordinates as well as intricate bio-chemical interactions between them that depend on their geometry. To be used in a learning pipeline, an initial modeling step transforming protein structures into a well-defined mathematical object is necessary. Different mathematical representations encode different structural and biological priors. For instance, the point cloud representation disregards the connectivity induced by chemical interactions, but allows for the most generic geometric description of the data. Protein surfaces choose to trade fine-grained information of the interior of a protein for an accurate depiction of its outer surface. This representation is thought to be of particular interest to study active interaction sites that mostly depend on properties of the surface because of the screening effect. However, even for interactions dominated by surface terms, the knowledge of the interior can encode the stability of the surface. These different representations are illustrated in Supplementary Figure˜1.
@@
Beyond using a single representation for proteins, the simultaneous representation of a protein as several mathematical objects holds promise. Indeed, different representations encode different biological priors of the data and present different computational advantages. A well-studied combination is the use of sequence information along with a graph representation of the structure. For instance, in Hermosilla et al. (2020); Fan et al. (2022) the authors enrich the graph with additional edge types that encode the sequence. Another way to include sequence information in a graph, is to use sequence embeddings, especially ones derived from protein language models and hence benefiting from the large amounts of available sequence data (Wu et al., 2023; Zhang et al., 2023). Finally, some approaches include information derived from protein structures in the training of protein language models (Bepler & Berger, 2019; Heinzinger et al., 2023; Su et al., 2023).
@@ 2 Motivation and Contribution
Despite this recent progress in comparing different representations for learning on protein data, relatively less focus has been given to surface-based representations, even though they have shown promising results in several applications (Gainza et al., 2020; Sverrisson et al., 2021; Wang et al., 2023). Approaches based on the surface representation have typically followed the initial MaSIF paper validation (Gainza et al., 2020), and hence have never been directly compared to other representations in the context of a single well-established benchmark. At the same time, powerful surface-based encoders have recently been proposed in the geometry processing/computer graphics literature, such as DiffusionNet (Sharp et al., 2022) significantly outperforming, in terms of both robustness and accuracy, the early Geodesic-CNN based techniques (Masci et al., 2015; Monti et al., 2017) which formed the basis of (Gainza et al., 2020). Unfortunately, it is not currently known how the best currently available surface-based encoders compare to other learning-based paradigms in the protein analysis tasks (e.g., on the Atom3d benchmark).
@@
  • State-of-the-art results in a wide variety of challenging scenarios and enhanced computational throughput by using residue graphs and coarsened meshes.
The rest of the paper is organized as follows: in Section˜3.1, we present the surface representation we use and the specialized networks used to process it. Section˜3.2 highlights a challenge for surface networks learning on various scales and presents solutions to mitigate these issues. In Section˜3.3, we propose to synergistically integrate graph and surface information within a unified architecture, harnessing the power of both representations. We provide the details regarding the chosen architecture in Section˜3.4 and analyze its computational aspects in Section˜3.5.
@@ 3.1 Surface Representation Learning
Our first objective is to study the utility of the best existing surface encoders for learning on protein data. To generate the surface representation $\mathcal{S}_{\mathbf{P}}$ of the protein $\mathbf{P}$, we rely on MSMS and on mesh coarsening and cleaning steps, detailed in Section˜C.1. At the basis for our surface-based learning method, we then employ the DiffusionNet approach (Sharp et al., 2022). This method has been proven to be highly robust and effective across a diverse set shape analysis tasks (e.g., (Attaiki et al., 2021; Cao & Bernard, 2022; Sun et al., 2023; Li et al., 2022) among others). In particular, DiffusionNet avoids using local patch parametrizations, which can lead to instabilities across different mesh structures and enables long range and multi-scale information propagation by using learned diffusion for information sharing. The mathematical foundation of DiffusionNet is the heat equation, which simulates the diffusion of heat, or, equivalently, the behavior of Brownian motion on a surface over time. For a surface $S$, let $\mathbf{f}_{t}:S\to\mathbb{R}$ be the function that defines the heat distribution over $S$ at time $t$ and $\Delta_{S}$ the Laplace-Beltrami operator (Meyer et al., 2003; Vallet & Levy, 2008) of the surface. The heat equation is the linear differential equation below (equation 1), solved by diagonalizing the Laplace-Beltrami. In practice, we store the $k=128$ smallest eigenvalues of $\Delta_{S}$ in a diagonal matrix $\Lambda\in\mathbb{R}^{k\times k}$, with the corresponding eigenvectors in $\Phi\in\mathbb{R}^{n\times k}$, and vertex area weights $M\in\mathbb{R}^{n\times n}$. The spectrally truncated solution to the heat equation is given as $\mathbf{f}_{t}=\Phi e^{-\Lambda t}(\Phi^{T}M)\mathbf{f}_{0}$.
$\text{{(Heat equation)} \quad\quad}\frac{\partial\mathbf{f}}{\partial t}$ $=\Delta_{S}\mathbf{f}.\phantom$$\text${{(Heat equation):} $\quad\quad$}} (1)
In DiffusionNet, this equation is used to perform information propagation on a surface using a feature map as $\mathbf{f}_{0}$ and learning the diffusion time in a task-specific manner. This mechanism relies on dense linear algebra operations, offering straightforward differentiation with respect to both $\mathbf{f}$ and $t$. The diffusion layers are combined with features based on $\mathbf{f}_{t}$, their spatial gradients and standard, point-wise MLPs. This leads to an architecture which can capture multi-scale geometric details of a surface in a task-specific manner.
@@ 3.2 Adapting to Protein Surfaces of Diverse Scales
To address this issue, we enhance the original DiffusionNet framework in two ways. First, we enable support for batch (the original model was limited to batch sizes of one) and incorporate a Batch Normalization layer (Ioffe & Szegedy, 2015) after each diffusion layer to stabilize learning. Second, we facilitate the optimization process by incorporating biological priors relevant to spatial scales. Consequently, we determined that diffusion times around 10 resulted in receptive fields around 10 Å (see Supplementary Figure˜2), which aligns with the spatial scale of binding sites. Inspired by the inherent multi-scale nature of protein structures, we opted to draw samples from a normal distribution $t\sim\mathcal{N}(10,5)$, characterized by a relatively high variance. The absolute values of these samples were then utilized as the initial values for our diffusion timescales. Both the large scale and the large variance are retained during training, as illustrated in Supplementary Figure˜3), enabling efficient multi-scale and long-distance message passing. These enhancements to DiffusionNet implementation mitigate the instabilities in the training process (as seen in Supplementary Figure˜3) and are available in the provided code repository and as a pip package.
@@ 3.3 Hybrid Representation Learning
As mentioned earlier, beyond assessing the efficacy of surface-based learning compared to other representations for protein learning, we explore the benefits of integrating different representations in a unified framework, harnessing their distinct strengths. Intuitively, surface representations can capture the intricate geometric details critical for tasks involving protein interactions, while graph representations detail the specific atomic interactions within a protein’s interior that indirectly influence its surface dynamics and interaction capabilities. Furthermore, these representations facilitate complementary approaches to learning: local message passing through graphs and global information dissemination for surfaces via learned diffusion. Inspired by these considerations, we propose a method that enables feature sharing between graph and surface-based representations. As emphasized below, unlike previous related approaches (Somnath et al., 2021; Pegoraro et al., 2024), we enable communication between the two representations across all learned layers of the network.
As a basis for our hybrid representation, in addition to the surface $\mathcal{S}_{\mathbf{P}}$, we construct a graph representation $\mathcal{G}_{\mathbf{P}}=(\mathcal{V}_{g},\mathcal{E}_{g})$. We use either a graph whose nodes are atoms, which aligns with the one used within the Atom3d benchmark, or one defined at the residue level. The residue-level graph is enriched with ESM-650M (Rives et al., 2021) sequence embeddings used as node features.
To construct our hybrid approach, we then build a bipartite graph $G=(V,E)$, where $V=\mathcal{V}_{g}\cup\mathcal{V}_{s}$ represents graph nodes and surface vertices, respectively. For each vertex on the surface, we find its 16 nearest neighbors in the graph and add the corresponding bidirectional edges in the bipartite graph. We provide a more detailed description of the construction and features of the atomic, residue and bipartite graphs in Section˜C.1.
@@
We now define block operations to encode a protein using $\mathcal{S}_{\mathbf{P}},\mathcal{G}_{\mathbf{P}}$ and $G$. Denote encoders on surfaces and graphs as $s_{\theta}$ and $g_{\theta}$, respectively, and the set of input features as $\mathcal{X}=\{x_{n},n\in\mathcal{V}\}$. The corresponding encoded features are $\mathcal{H}=\{h_{n},n\in\mathcal{V}\}$ with $h_{n}=s_{\theta}(x_{n})$ for nodes $n\in\mathcal{V}s$ and $h_{n}=g_{\theta}(x_{n})$ for nodes $n\in\mathcal{V}_{g}$. Our general methodology incorporates message-passing neural networks, denoted $\text{MP}_{\theta}$, over the bipartite graph $G$, such that at a layer $l$, we get $\mathcal{X}^{l+1}=\text{MP}^{l}_{\theta}(\mathcal{H}^{l})$. By employing distinct sets, $\theta_{sg}$ and $\theta_{gs}$, the architecture handles messages traversing from the surface to the graph and vice versa. Those block operations can be stacked as shown in Figure˜1. We emphasize that our feature sharing occurs on the local (node) level and is enabled by the proximity relations in 3D space. Moreover, our hybrid approach trains both representations jointly, while enabling information sharing across all network layers, which is crucial to its success.
@@ 3.4 Proposed architectures
Different bipartite message passing networks and organization are possible, encompassing as special cases several existing approaches. We provide the implementation details in Appendix˜C. We evaluate these configurations and conduct several ablation studies on our final model, in Section˜4.5.
@@ 3.5 Computational enhancements
Surface-methods are traditionally thought to be compute-expensive methods, motivating approaches to side-step their intrinsic complexity (Sverrisson et al., 2021). Through a complexity analysis presented in Section˜D.1, we found the number of vertices to be critical in DiffusionNet runtime, which we addressed by coarsening our meshes. In this coarse surface regimen, the graph encoding with ProNet is the computational bottleneck, which is smaller than when used in isolation. Hence, we claim that our method is faster than this efficient graph encoder in the coarse surface regimen.
From the storage and memory perspectives however, the memory footprint of the surface-related operations dominate the ones originating from the graphs, especially because of stochasticity. We did not find it to be limiting in terms of I/O, but rather in terms of batch size. We implemented a dynamic batching procedure, alleviating our memory issues. Further details are provided in Section˜D.2. Finding ways to reduce the memory footprint of surface networks remains an important direction.
@@ 4.1 Performance of Surface Representation
We start by validating our surface encoder on the RNA segmentation benchmark (Poulenard et al., 2019) for surface methods. The task is to segment 5s ribosomal RNA molecules into functional components. This dataset consists of 640 RNA surface meshes of about 15k vertices, and was already used to compare modern surface encoders. We assess the impact of the proposed enhancements to DiffusionNet by showing the learning curves of the enhanced models on the RNA segmentation task (see Figure˜3 and Section˜F.1 for a similar analysis on PSR). Moreover, we compare their performance to other recent surface encoders, DGCNN (Wang et al., 2019) and DeltaConv (Wiersma et al., 2022) and report results in Figure˜3. Our experiments show that our adjustments significantly improve the DiffusionNet stability issues, and allow it to converge to a model with a test accuracy of 84.1% instead of of 80.9%. This accuracy is significantly higher than DGCNN (74.7%) and DeltaConv (78%).
[…]
Then, to compare the performance of our surface encoder to the use of other representations, we turn to the Atom3d benchmark, focusing on its three tasks exclusive to proteins. The Protein Interaction Prediction (PIP) task aims to predict which part of a protein interacts with another, and holds about 120k examples. The Mutation Stability Prediction (MSP) task is to determine if a mutation enhances the stability of protein-protein interaction (5k examples). Finally, the Protein Structure Ranking (PSR) task aims to assign a quality score to predicted protein structures with around 10k proteins.
A more detailed description of all our datasets is available in Appendix˜E.
@@
Auroc Auroc $R_{l}$ $R_{g}$